OpenStudy (anonymous):

Which of the following represents the zeros of f(x) = 4x3 − 13x2 − 37x + 10? what i think x = 5 x = -2 x = 1/4 @plasmafuzer

OpenStudy (anonymous):

these are the other answers and stuff

OpenStudy (anonymous):

i just need reassurance like is it the right one or not

OpenStudy (anonymous):

Yea your answers are good

OpenStudy (anonymous):

When you finish come back to this post since it isn't as cluttered... and we will work back through the last problem.

OpenStudy (anonymous):

perfect!! thanks, i will

OpenStudy (anonymous):

Awesome, also may I have a medal plz :D

OpenStudy (anonymous):

For reference: (x-4)(x+5)(x-7)

OpenStudy (anonymous):

well deserved C:

OpenStudy (anonymous):

Cheers :D

OpenStudy (anonymous):

I got 100

OpenStudy (anonymous):

CONGRATS!!!

OpenStudy (anonymous):

Do you still want to help me with the other one? Or would you like to do it later, sorry i feel like I'm taking up your time

OpenStudy (anonymous):

Ok now comes the real taskmaster me hehehe

OpenStudy (anonymous):

Nope your not getting out of it that easily... Im alright Im relaxing with a keyboard on a nice Saturday

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

perfect, okay so explain this to me please! How did you get the answer

OpenStudy (anonymous):

Ok so given (x-4)x distribute this for me plz

OpenStudy (anonymous):

Ill walk you through one piece at a time and once we finish I will post a totally start to finish polished solution ok?

OpenStudy (anonymous):

But YOU will do the work each step I will just make sure you do it right.

OpenStudy (anonymous):

So please: (x-4)x = ?

OpenStudy (anonymous):

I sense uneasiness whats up?

OpenStudy (anonymous):

Please tell me whenever you are confused.

OpenStudy (anonymous):

okay sorry my page froze and i was trying to refresh (mess) (x-4)x x*4-x*x

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Umm yes... except you somehow swtiched the minus sign.

OpenStudy (anonymous):

Try it again please.

OpenStudy (anonymous):

a(b – c) = a × b – a × c I'm going by this, ok lets try it again? (x-4)x x*x-4*x

OpenStudy (anonymous):

There we go brilliant :D

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

And is there another way we can write x*x?

OpenStudy (anonymous):

x^2?

OpenStudy (anonymous):

BRILLIANT!!!!

OpenStudy (anonymous):

Ok so in other words (x-4)x=x^2-4x yes? what about: (x-4)5

OpenStudy (anonymous):

Note the multiplication is implied in a statement written like this

OpenStudy (anonymous):

(x-4)5 x*5-4*5 (4 has to automatically go next to the minus sign because its actually -4 thats why id messed up earlier right, i didn't catch that)

OpenStudy (anonymous):

YES EXACTLEY :D... Now simplfy it completely

OpenStudy (anonymous):

i.e. get rid of the * signs

OpenStudy (anonymous):

ok, 4*5=20 but what would i do with x*5?

OpenStudy (anonymous):

what does x*5 mean write it out in words

OpenStudy (anonymous):

some number multiplied by five

OpenStudy (anonymous):

is that the same as 5 times some number?

OpenStudy (anonymous):

basically yeah

OpenStudy (anonymous):

basically nothing that is an important point... you may not recall but way back when you were young your learnd multiplication is communicative... that means for ALL... let me say this again FOR ALL real (and later on complex) numbers a*b=b*a

OpenStudy (anonymous):

yes thats rights, doesn't matter how you put it its still the same thing

OpenStudy (anonymous):

so saying x*5=5*x is a fundamental point about basic arithmetic its called a field axiom... without it math wouldnt work... think about that

OpenStudy (anonymous):

like 5*x= x*5

OpenStudy (anonymous):

EXACTLEY :D :D

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

but how would i simplify this to get rid of the * and finish the equation?

OpenStudy (anonymous):

Ok so it happens to be the convention in algebra that when you have a number times a variable (some number x) you put the number first... and you can just drop the multiplication sign.... it is implied

OpenStudy (anonymous):

Bingo

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

So what is (x-4)5=

OpenStudy (anonymous):

is that it? 5x-20

OpenStudy (anonymous):

Boooomm like you have been doing it for years :D

OpenStudy (anonymous):

Ok lets put it together now... (x-4)(x+5)

OpenStudy (anonymous):

really????!! that was it? i was expecting something super difficult (':

OpenStudy (anonymous):

Everyone always does when it comes to math... it comes down to some fundamental uncertainties they have about some basic maths

OpenStudy (anonymous):

Ill tell you my story when we finish

OpenStudy (anonymous):

(x-4)5 x*5-4*5 x*5= 5x 4*5=20 5x-20

OpenStudy (anonymous):

Ok what about (x-4)(x+5)???

OpenStudy (anonymous):

Note you have basically solved it already... just in two parts...work it out step by step please and post each step as you go

OpenStudy (anonymous):

Dont be afraid... take the leap

OpenStudy (anonymous):

what if it was a(x+5) what would the answer be?

OpenStudy (anonymous):

a(x+5)=?

OpenStudy (anonymous):

@emilybbydoll Did I lose you?

OpenStudy (anonymous):

sorry i was v hungry i had to go get a snack

OpenStudy (anonymous):

im lost on how to (x-4)(x+5) i could the other one v good but idk about this

OpenStudy (anonymous):

@plasmafuzer

OpenStudy (anonymous):

Ok what about: a(x+5)

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

Dont overthink... just do what you know you need to

OpenStudy (anonymous):

a(x+5)

OpenStudy (anonymous):

a(x+5) x*a+5*a

OpenStudy (anonymous):

Bingo... and since a is an unknown do we need the *'s?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we didnt with the x's earlier... and it is just an unknown

OpenStudy (anonymous):

so what do you think now?

OpenStudy (anonymous):

does it matter what I call something?

OpenStudy (anonymous):

but how would i go about it? what do i distribute first? thats what I'm confused about

OpenStudy (anonymous):

relax dont overthink.... all we have is a(x+5) right now so lets deal with that

OpenStudy (anonymous):

you correctly distributed it but the question was about the *'s

OpenStudy (anonymous):

does it matter what i name a variable?

OpenStudy (anonymous):

x(a)+5a ? oh nvm i was talking about (x-4)(x+5)

OpenStudy (anonymous):

we arent there yet... ok now you have parentheses... did we need them when we had x(x)?

OpenStudy (anonymous):

you could've used it but i guess its preferable to use x^2

OpenStudy (anonymous):

Good point... we didn't write it that was because x^2 is the BETTER way, but what about 5(x)?

OpenStudy (anonymous):

at the end of the day x just stands in for some number and you were perfectly comfortable writing 5x yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if a is just standing in for something why not just ax?

OpenStudy (anonymous):

yes... no?

OpenStudy (anonymous):

it doesn't sound right?

OpenStudy (anonymous):

Ok ok :) tell me why.. I just told you why it does sound right tell me why it doesnt.

OpenStudy (anonymous):

Forget what I may or may not know.... If x is a variable that stands in for something then I can write 5x no problem you agree... but why cant I write ax is a just stands in for something... what if it was 5.. a=5 i mean

OpenStudy (anonymous):

im guessing its informal, it just sounds like your writing something tbh

OpenStudy (anonymous):

No not at all this is totally formal and totally allowed... my freely defining things the way that I am (as you know dont already know) is due to my ability freeing me up to be creative with the "language" that is math.